OpenStudy (kainui):

I felt like making up an integral for people to try to solve for fun. Good luck!

OpenStudy (kainui):

OH NO! I had a typo in my question, the 9 in the numerator was supposed to be a 4! \[\int\limits_{-9}^{9}\frac{ (36-4x^2)^{3/2} }{ 9-x^2 }dx\]

zepdrix (zepdrix):

oh :c

OpenStudy (kainui):

Sorry Zepdrix. Hopefully that makes it easier. >_> I'm so sorry, I checked it like 10 times and typed the wrong thing ugh lol.

zepdrix (zepdrix):

That makes it quite it bit easier to work through c: heh

OpenStudy (kainui):

I'm just experimenting with creating problems, since I'm usually just solving them. If you're interested in making some or if you've seen some particularly fun/interesting/tricky ones please feel free to share lol.

OpenStudy (kainui):

Well, promise this is the last correction, the limits of integration are -3 to 3, I think I'm tired from too many finals and studying all week... I swear it's an entirely legitimate problem this time.

zepdrix (zepdrix):

Mmmmmmmmmmmmm I came up with 96ish.. am i close am i close??? O:

OpenStudy (kainui):

I might as well write out the steps and go to sleep, but you are rather close. I'm curious what you did, the answer should contain pi in it.

zepdrix (zepdrix):

Hmmm. I did a trig sub, x=3sinQ, dx=3cosQ dQ. There were quite a few steps :) It's possible I goofed up somewhere heh

OpenStudy (kainui):

\[4\int\limits_{-3}^{3}\frac{ (4(9-x^2))^{3/2} }{ 4(9-x^2) }dx\]\[4\int\limits_{-3}^{3}\sqrt{4(9-x^2)}dx\]\[8\int\limits_{-3}^{3}\sqrt{9-x^2}dx\]From here you can see that this is a semicircle with radius of 3, so the integral becomes\[8*\frac{ \pi 3^2 }{ 2 }=36\pi \approx 113.0973355\]

zepdrix (zepdrix):

Oh neato :o

OpenStudy (kainui):

Haha I'm scared to imagine what trig substitution would bring! I almost want to try it and find out, but yeah I saw an integral like this today and was pretty shocked and looks awesome lol. =D

OpenStudy (anonymous):

 subbing function: \[\large \sin\theta = \frac{1}{3}\sqrt{9-x^2}\] \[\large 3 \sin \theta = \sqrt{9-x^2}\] subbing derivative: \[\large \cos\theta=\frac{1}{3}x\] \[\large x =3\cos\theta\] \[dx=-3\sin\theta \space d\theta\]